# supplement_jasa

### v2

@@ 2.3 Accelerated learning for spatial random fields
where $\nu^{\star}_{[-l]}=\nu_{0}+n_{k,[-l]}$, $n_{k,[-l]}$ is the cardinality of $\mathcal{S}_{k,[-l]}$ (which is the set of locations in $\mathscr{D}_{k,[-l]}$), and $\Psi^{\star}_{[-l]}=\Psi_{0}+(\alpha_{j}^{-1}-1)Y^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu^{{\mathrm{\scriptscriptstyle T}}\star}_{\gamma,[-l]}V^{-1\star}_{[-l]}\mu^{\star}_{\gamma,[-l]}$. While $V^{\star}_{i}=M_{y,i}V_{\gamma}^{\star}M_{y,i}^{{\mathrm{\scriptscriptstyle T}}}+V_{\Omega_{i}}+(\alpha_{j}^{-1}-1)$, $\mu^{\star}_{i}=M_{y,i}\mu^{\star}_{\gamma,[-l]}$, which are defined by computing the following auxiliary quantities: $V^{-1\star}_{\gamma,[-l]}=\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}X_{k,[-l]}+M_{0}^{-1}&\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}\\ \frac{\alpha_{j}}{1-\alpha_{j}}X_{k,[-l]}&\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})+\frac{\alpha_{j}}{1-\alpha_{j}}\mathbb{I}_{n_{k,[-l]}}\end{bmatrix}$, $\mu^{\star}_{\gamma,[-l]}=V^{\star}_{\gamma,[-l]}\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}\\ \frac{\alpha_{j}}{1-\alpha_{j}}Y_{k,[-l]}\end{bmatrix}$, $M_{y,i}=\begin{bmatrix}X_{k,i}&\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\end{bmatrix}$, and $V_{\Omega_{i}}=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,i})-\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\rho_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,i})$. Note that $\nu^{\star}_{[-l]}$ is scalar, $\mu^{\star}_{i}$ is $(1\times q)$ row vector, $V^{\star}_{i}$ is a scalar, and $\Psi^{\star}_{[-l]}$ is $(q\times q)$ matrix. Further details, including derivations and implementation, are supplied in Section LABEL:sec:matrixT and Algorithm 1.
@@
For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\cdot\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$. Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\cdot\mid\mathscr{D}_{k})$ over the $k$ subsets. This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{i}\mid\mathscr{D}_{k}\right)$ (see Section LABEL:sec:BPSobjfun). Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
$\hat{p}(\cdot\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)\;.$ (2.10)
@@
First, the set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$ obtained using bps within the subset of the data $\mathscr{D}_{k}$, is primarily used to approximate the subset posterior distribution $\hat{p}(\Theta\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}p\left(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$, for each subset $k=1,\dots,K$. By considering the second set of stacking weights $\hat{w}=\{\hat{w_{k}}\}_{k=1,\dots,K}$, the stacked full posterior distribution is $\hat{p}(\Theta\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Theta\mid\mathscr{D}_{k})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j})$, where $\Theta=\{\gamma,\Sigma\}$. This is a substantial simplification over meta-kriging (Guhaniyogi and Banerjee,, 2018, 2019) that does not require any empirical approximation of posterior or predictive distributions. Again, the predictive random variable $Y_{\mathcal{U}}$ is recovered from (2.10) as
$\hat{p}(Y_{\mathcal{U}}\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(Y_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right).$ (2.11)
Inferential interest resides with the posterior predictive surface for $\Omega$. Thus, we estimate $\Omega_{\mathcal{U}}$ using $\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k})$, where $\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$.
@@
3 Computer programs and resources
Our subsequent implementations, including all simulations experiments and data analysis, are executed in native R and c++ deploying the spBPS package. In particular, all programs used to obtain the ensuing analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI that links the Rcpp-based spBPS package including the main function executing bps for univariate and multivariate geostatistical models. The reported results are executed on a standard laptop, which runs an Intel Core I7-$8750$H CPU with $5$ available cores for parallel computation and $16$ GB of RAM.
We conduct experiments comparing double bps with spatial meta-kriging (smk) using a meta-posterior (Minsker et al.,, 2017) defined as the geometric median of individual posterior distributions based on subsets of the data (as implemented in Guhaniyogi and Banerjee,, 2018, 2019). We execute meta-kriging using the spMvLM function within the spBayes R package that fits a linear model of coregionalization for multivariate spatial data using MCMC (Finley et al.,, 2015). For parallel implementations of bps and smk, we employ R packages doParallel, and foreach (Corporation and Weston,, 2022; Microsoft and Weston,, 2022). We map the interpolated spatial surfaces using the MBA R package (Finley et al.,, 2022), while sampling from the matrix-variate normal and t distributions are achieved using the mvnfast R package (Fasiolo,, 2014). Section LABEL:sec:compdetails in the Supplement provides further details including memory management considerations and efficient solutions to spatial “BIG” data analysis, and sensitivity on the number of data shards $K$.
@@ 4 Simulation experiments
We conduct multiple simulation experiments to evaluate inferential performance of double bps while underscoring comparisons with smk. We present a selection of simulation results for multivariate models here and refer the reader to Section LABEL:sec:sim_suppl for further experiments.
4.1 Simulation - BPS vs SMK
In the first setting, we consider two synthetic datasets with common structures but distinct sizes using the model (2.4). Both datasets consist of $p=2$ predictors and $q=2$ response variables, but different numbers of spatial locations, $n=5,000$ and $n=10,000$, respectively. We generate spatial coordinates from a uniform distribution on $[0,1]^{2}$ domain and fix them. We build the $n\times n$ spatial correlation matrix $V$ over these coordinates using an exponential spatial correlation function with $\phi=4$ and specify $\Sigma=\mathbb{I}_{q}$. From these specifications we generate the $n\times q$ matrix $Y$ from the first equation of (2.4) with fixed $p\times q$ matrix $\beta=\begin{bmatrix}-0.75&1.85\\ 0.90&-1.10\end{bmatrix}$, a fixed $n\times p$ matrix $X$ with a first column of ones, representing the intercept, and $p-1$ columns of values randomly simulated from a uniform distribution on $[0,1]$ (emulating standardized predictors), and the proportion of spatial variability $\alpha=0.8$.
We analyze the simulated datasets using double bps and smk in two settings: (i) $K=10$ and $n=5,000$; (ii) $K=10$ and $n=10,000$; (iii) $K=5$ and $n=5,000$; and (iv) $K=20$ and $n=10,000$. These settings produce subsets of size $n/K\in\{500,1000\}$. We implement bps using $J=9$ candidate models $\mathscr{M}_{j},j=1,\dots,J$, where each model is specified by a set of candidate values for the hyperparameters ${\alpha_{j},\phi_{j}}$ in (2.4). These hyperparameters represent the proportion of spatial variability and the parameter(s) of the spatial correlation function, respectively. The set of candidate models is constructed as the set of all possible combinations of values for these hyperparameters. In the subsequent experiments, the grid of models was built using $\alpha\in\{0.70,0.80,0.90\}$ and $\phi\in\{3,4,5\}$. These values resemble an effective spatial range of $\{0.99,0.75,0.60\}$ units, corresponding to $70\%,53\%,42\%$ of the maximum inter-site distance inside the unit square, beyond which the spatial correlation drops below $0.05$. Equation (2.5) follows choices in Zhang et al., (2021). Specifically, we set $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, and $\nu_{0}=3$ in the $\operatorname{MNIW}$ joint prior for $\{\gamma,\Sigma\}$ in 2.5. For data analysis, we use an exponential spatial correlation function for $\rho_{\phi}(\cdot,\cdot)$, which is completely defined conditionally on $\mathscr{M}_{j}$, since specifies a value for $\phi$. Again, in the conjugate framework, we draw $R=250$ posterior samples used for inference.
Model $n=5000,K=10$ $n=5000,K=5$ $n=10000,K=20$ $n=10000,K=10$
smk 51.41 237.23 103.36 446.01
bps 1.38 7.22 2.24 10.58
Relative ratio 37.25 32.85 46.14 42.15
Table 1: Computational aspects for the different settings, with time in minutes.
We also apply smk to the two simulated datasets with the same combinations of $n$ and $K$ as bps. Unlike bps, where we stack analytically tractable posteriors over a range of fixed values of spatial covariance kernel parameters, the smk implementation attempts full Bayesian inference using prior distributions on spatial covariance kernel parameters. We fit the linear model of coregionalization described in Finley et al., (2015) for each subset of the multivariate spatial data using MCMC. The posterior samples from the $K$ subsets are combined using Weiszfeld’s iterative algorithm (Minsker et al.,, 2017) to produce an estimate of the geometric median of the posterior distributions.
Table 1, which compares the computational speed of double bps with smk reveals massive computational gains accrued from bps. The computational advantage evinced from the relative ratio become more pronounced as the size of the data becomes larger despite the larger subsets. This is explained by the fact that fitting the Gaussian process regression dominates the computation relative to the assimilation of inference from the subsets. If the number of locations explode, then the geometric median of posteriors required by smk is computationally unfeasible. While smk offers Bayesian estimates using MCMC for each subset, double bps avoids MCMC and, hence, issues of convergence.
Importantly, as we see in Figure 2, which depicts estimated response surfaces using bps and smk corresponding to $n=5000$ and $K=10$, inference from smk and bps are practically indistinguishable. The root mean squared prediction errors, rmspe, reported in Figure 2 denotes the average squared differences between the generated and estimated values of the response and reveals minor discrepancies between bps and smk. Section LABEL:sec:sim_suppl_M presents results for the other configurations of $n$ and $K$, each of which reveals that bps offers practically indistinguishable spatial interpolation from smk at a fraction of the computational cost.
Figure 2: from left to right: comparison between the true generated response surfaces, the surfaces predicted from bps and smk (posterior mean), with rmspe. For $n=5000$, $K=10$.
Figure 3 reports 95% posterior predictive intervals for the response. Again, the empirical coverage is impressive. While we see slightly wider intervals from BPS, this is less pronounced than the underestimation of variability seen with smk. Moreover, Figure 3 reveals superior map estimates for the bps. Finally, Figure 4 presents the recovery of parameter estimates. As seen in predictive inference, the posterior credible intervals for parameters also deliver practically indistinguishable inference for the two modeling frameworks. In particular, both methods recover the true values for $\beta$ and $\Sigma$, while BPS reconstructs a better point estimate for range parameters $\phi$ using $\sum_{k=1}^{K}\hat{w}_{k}\sum_{j=1}^{J}\hat{z}_{j}\phi_{j}$. Similar conclusions are drawn from other values of $n$ and $K$ explored in Section LABEL:sec:sim_suppl for univariate responses.
Figure 3: from top to bottom: comparison between posterior predictive intervals for the predicted response from bps and smk, with empirical coverage. For $n=5000$, $K=10$.
Figure 4: from left to right: comparison between posterior credible intervals for the parameters recovered from bps and smk. For $n=5000$, $K=10$.
4.2 Simulation - Multivariate Transfer learning
We now consider inferential performance for different misspecified models. As in the earlier example, we generate the $n\times q$ outcome $Y$ using  (2.4) with $n=5,000$, $q=3$ and $p=2$, $X$ includes an intercept and $p-1$ columns generated from a standard uniform distribution over $[0,1]$, $\beta=\begin{bmatrix}-0.75&1.05&-0.35\\ 2.20&-1.10&0.45\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2.00&0.80&0.20\\ 0.80&2.00&-0.45\\ 0.20&-0.45&2.00\end{bmatrix}$. The $n\times n$ spatial correlation matrix $V$ is specified using an exponential correlation matrix with a fixed range parameter at $\phi=4$ and $\alpha=0.8$ to be the strength of spatial dependence.
We analyze the generated data using three different models that assign prior distributions to $\{\beta,\Sigma\}$ with $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, and $\nu_{0}=3$, but different values of $\alpha$ and $\phi$ representing (i) well-specified setting with the data generating values $\{\alpha=0.8,\phi=4\}$; (ii) moderately misspecified setting with $\{\alpha=0.45,\phi=6.63\}$; and (iii) a highly misspecified setting with $\{\alpha=0.25,\phi=50\}$. For each setting, we partition the original data into $K=10$ subsets, corresponding to subsets of size $500$. We consider $J=9$ competitive models for bps and consider $\alpha\in\{0.75,0.80,0.85\}$ and $\phi\in\{2,4,6\}$ that yields effective spatial ranges in the percentage of maximum point inter-distance of $105\%,53\%,35\%$ respectively.
Here, we display results for the bps, while graphical results for other settings are available in Section LABEL:sec:sim_suppl_M. Figure 5 presents a visual comparison between the generated true response and the posterior mean of spatial predictions delivered by bps. We present the rmspe along with a set of credible intervals for the parameters relative to the true values. The inferential performance evaluated using the rmspe as well as the spatially interpolated maps corroborates the effectiveness of bps. Figure 6 reveals that more than half of the model unknowns are included in the 95% posterior credible interval. Unsurprisingly, inferential performance is weakest for the spatial variance, which is not identifiable from the realized data (Zhang,, 2004). Further results on simulation experiments are available in Section LABEL:sec:sim_suppl_M.
Figure 5: from top to bottom: comparison between the true generated spatial surfaces and the posterior mean surfaces of predictions from bps, with rmspe for each outcome.

### v3

@@ 2.3 Accelerated learning for spatial random fields
where $\nu^{\star}_{[-l]}=\nu_{0}+n_{k,[-l]}$, $n_{k,[-l]}$ is the cardinality of $\mathcal{S}_{k,[-l]}$ (which is the set of locations in $\mathscr{D}_{k,[-l]}$), and $\Psi^{\star}_{[-l]}=\Psi_{0}+(\alpha_{j}^{-1}-1)Y^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu^{{\mathrm{\scriptscriptstyle T}}\star}_{\gamma,[-l]}V^{-1\star}_{[-l]}\mu^{\star}_{\gamma,[-l]}$. While $V^{\star}_{i}=M_{y,i}V_{\gamma}^{\star}M_{y,i}^{{\mathrm{\scriptscriptstyle T}}}+V_{\Omega_{i}}+(\alpha_{j}^{-1}-1)$, $\mu^{\star}_{i}=M_{y,i}\mu^{\star}_{\gamma,[-l]}$, which are defined by computing the following auxiliary quantities: $V^{-1\star}_{\gamma,[-l]}=\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}X_{k,[-l]}+M_{0}^{-1}&\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}\\ \frac{\alpha_{j}}{1-\alpha_{j}}X_{k,[-l]}&\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})+\frac{\alpha_{j}}{1-\alpha_{j}}\mathbb{I}_{n_{k,[-l]}}\end{bmatrix}$, $\mu^{\star}_{\gamma,[-l]}=V^{\star}_{\gamma,[-l]}\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}\\ \frac{\alpha_{j}}{1-\alpha_{j}}Y_{k,[-l]}\end{bmatrix}$, $M_{y,i}=\begin{bmatrix}X_{k,i}&\quad\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\end{bmatrix}$, and $V_{\Omega_{i}}=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,i})-\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\rho_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,i})$. Note that $\nu^{\star}_{[-l]}$ is scalar, $\mu^{\star}_{i}$ is $(1\times q)$ row vector, $V^{\star}_{i}$ is a scalar, and $\Psi^{\star}_{[-l]}$ is $(q\times q)$ matrix. Further details, including derivations and implementation, are supplied in Section LABEL:sec:matrixT and Algorithm 1.
@@
For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\,\cdot\;;\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\,\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$.Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\,\cdot\;;\mathscr{D}_{k})$ over the $k$ subsets. This double bps (dbps) of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{i}\mid\mathscr{D}_{k}\right)$ (see Section LABEL:sec:BPSobjfun). Once stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are computed using Algorithm 2, sampling from the posterior and posterior predictive distributions is obtained from
$\hat{p}(\,\cdot\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\,\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)\;.$ (2.10)
@@
First, the set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$ obtained using bps within the subset of the data $\mathscr{D}_{k}$, is primarily used to approximate the subset posterior distribution $\hat{p}(\Theta\;;\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}p\left(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$, for each subset $k=1,\dots,K$. By considering the second set of stacking weights $\hat{w}=\{\hat{w_{k}}\}_{k=1,\dots,K}$, the stacked full posterior distribution is $\hat{p}(\Theta\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Theta\;;\mathscr{D}_{k})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j})$, where $\Theta=\{\gamma,\Sigma\}$. This is a substantial simplification over meta-kriging (Guhaniyogi and Banerjee, , 2018, 2019) that does not require any empirical approximation of posterior or predictive distributions. Again, the predictive random variable $Y_{\mathcal{U}}$ is recovered from (2.10) as
$\hat{p}(Y_{\mathcal{U}}\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(Y_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right).$ (2.11)
Inferential interest resides with the posterior predictive surface for $\Omega$. Thus, we estimate $\Omega_{\mathcal{U}}$ using $\hat{p}(\Omega_{\mathcal{U}}\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Omega_{\mathcal{U}}\;;\mathscr{D}_{k})$, where $\hat{p}(\Omega_{\mathcal{U}}\;;\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$.
@@
3 Computer programs and resources
All our subsequent analysis are implemented in native R and c++ deploying the spBPS package. All programs required to reproduce the analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI that links the Rcpp-based spBPS package. The reported results are from a standard laptop running an Intel Core I7-87508750H CPU with $5$ cores for parallel computation and $16$ GB of RAM.
We fit a linear model of coregionalization (lmc, Banerjee et al., , 2015) and multivariate seemingly unrelated Bayesian additive regression trees (multivariate bart Esser et al., , 2025) using spBayes and suBART packages, respectively. We also compare with machine learning methods and AI systems using a scalable platform for parallelized supervised and unsupervised machine learning algorithms offered by h2o (Fryda et al., , 2024). We specifically fit distributed random forest (drf), gradient boosting (gbm), deep neural network (dnn), and a fully automatic machine learning algorithm (automl). For parallel implementations of dbps, we employ R packages doParallel, and foreach (Corporation and Weston, , 2022; Microsoft and Weston, , 2022). We map the interpolated spatial surfaces using MBA (Finley et al., , 2022), while sampling from the matrix-variate normal and t distributions is achieved using mvnfast (Fasiolo, , 2014). Section LABEL:sec:compdetails specifies computational considerations and sensitivity to the number of data shards, $K$, for spatial “BIG” data analysis.
We build a Bayesian transfer learning engine to conduct amortized Bayesian inference (Zammit-Mangion et al., , 2024) using dbps. We implement a residual neural network (ResNet) (He et al., , 2015) using the R interfaces supplied by tensorflow (Allaire et al., , 2024) and keras (Kalinowski et al., , 2024) for native Python.
@@ 4 Simulation experiments
We evaluate computational and inferential performances of double bps, while underscoring comparisons with multiple alternative methodologies. We present a selection of simulation results for multivariate models here and refer the reader to Section LABEL:sec:sim_suppl_M for further experiments.
4.1 Simulation - Transfer learning in $\mathscr{M}$-closed & $\mathscr{M}$-open settings
We evaluate inferential and predictive performance under different settings. We explore how double bps behaves in the $\mathscr{M}$-closed and $\mathscr{M}$-open settings and compare with an exact transfer learning framework we devise in Section 2.1. In the latter, model specification is characterized by different values of $\alpha$ and $\phi$ representing (i) well-specified (ws) setting with the data generating values $\{\alpha=0.8,\phi=4\}$; (ii) moderately misspecified (ms) setting with $\{\alpha=0.45,\phi=6.63\}$; and (iii) a highly misspecified (hms) setting with $\{\alpha=0.25,\phi=50\}$.
The double bps was tested over $\mathscr{M}$-closed and $\mathscr{M}$-open settings. The former considers situations where the true model exists, and it is identified within a finite set of considered models. Here, the “true” model is the one such that $\{\alpha=0.8,\phi=4\}$. Then, for double bps under closed setting (bps-c) we specify $J=9$ competitive models with $\alpha\in\{0.75,0.80,0.85\}$ and $\phi\in\{2,4,6\}$ that yields effective spatial ranges in the percentage of maximum point inter-distance of $105\%,53\%,35\%$ respectively, including the true model as one of the possible candidates. Conversely, in the $\mathscr{M}$-closed setting, even though the true model exists, it cannot be fully specified. Thus, for double bps under open setting (bps-o), we randomly define $J=9$ candidate models. In particular, we uniformly sample 3 values for $\alpha\in(0,1)$ and 3 values for $\phi\in(0,50)$. We perform the experiment using 50 replications. Each replicate consists of values of $n\times q$ outcome $Y$ generated from (2.4) with $n=5,000$, $q=3$ and $p=2$, $X$ includes an intercept and a predictor generated from a standard uniform distribution over $[0,1]$, $\beta=\begin{bmatrix}-0.75&1.05&-0.35\\ 2.20&-1.10&0.45\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2.00&0.80&0.20\\ 0.80&2.00&-0.45\\ 0.20&-0.45&2.00\end{bmatrix}$. The $n\times n$ spatial correlation matrix $V$ is specified using an exponential correlation function with $\phi=4$ and $\alpha=0.8$.
Figure 2 presents (i) mean square prediction error (mspe); (ii) predictive interval width; (iii) absolute bias; and (iv) variance. We present boxplots for the distribution of each metric over the 50 replicates. This is made for each response and for each setting. In terms of predictive mspe, absolute bias, and variance, the settings bps-c, bps-o depict sligthly better performance. However, this seems to be compromising predictive interval width as the uncertainty of prediction results is much higher. Approaches that estimate $\{\alpha,\phi\}$, rather than fix them, more uncertainty is somehow expected. In addition, we find no evidence of any difference between bps-c, and bps-o for any metric. This is surprising, as it suggests the reliability of dbps even under the transfer learning setting we devise in Section 2.3. Finally, irrespective of $\mathscr{M}$-closed or $\mathscr{M}$-open settings, it is more convenient to specify a set of candidate models using double bps instead of attempting to fix $\{\alpha,\phi\}$.
Figure 3 presents posterior inference for (i) average empirical bias; (ii) average coverage; and (iii) average standard deviation, where the average is taken over the 50 replications. As expected, misspecification induces empirical bias in posterior estimates. The top panel reports how greater levels of bias are associated with models that are farther away from the truth. Here, double bps is placed in the middle for both $\mathscr{M}$ settings. The middle plot in Figure 3 shows that $\Sigma$ is the most affected parameter by misspecifications. Its elements are well captured only by double bps, with coverage close to nominal, followed by ws specification, which performs worse. A Similar pattern holds for $\beta$, where only double bps ensures adequate coverage. This reflects greater posterior variability in stacking approaches, as shown in the bottom panel reporting posterior standard deviations across settings. Unsurprisingly, inferential performance is weakest for the spatial variance, which is not identifiable from the realized data (Zhang, , 2004). double bps behaves very similarly among $\mathscr{M}$-closed and $\mathscr{M}$-open. Additional simulation experiments are discussed in Section LABEL:sec:sim_suppl_M.
Figure 2: Predictive MSPE, interval width, absolute bias, and variance boxplot across responses and settings from 50 replications.
Figure 3: Average posterior bias, coverage, and standard deviation across parameters and settings from 50 replications.
While dbps outperforms exact transfer learning (Section 2) in $\mathscr{M}$-closed and $\mathscr{M}$-open settings, the latter performs competitively with improved predictive performance in ws and ms settings than with hms. Our overall findings appear consistent with theoretical insights that Gaussian processes tend to deliver good predictive performance even for misspecified covariance functions in fixed domains (Stein, , 1988; Stein and Handcock, , 1989).
4.2 Amortized Bayesian inference
We conduct transfer learning by supervising a neural network using the outputs of double bps to deliver amortized Bayesian inference. We generate $100$ instances of $Y$ from (2.4) using a fixed realization of $\Omega$ for $q=2$ correlated outcomes, $n=500$ spatial locations that remain fixed across the datasets, and a fixed design matrix $X$ with $p=2$ comprising an intercept and a single predictor whose values were sampled independently from a uniform distribution over $[0,1]$. The true regression coefficients are fixed at $\beta=\begin{bmatrix}-0.75&1.85\\ 0.9&-1.10\end{bmatrix}$, with $\Sigma=\begin{bmatrix}1&-0.3\\ -0.3&1\end{bmatrix}$, $\alpha=0.8$, and $\rho_{\phi}(s_{i},s_{j})=\exp(-\phi\|s_{i}-s_{j}\|)$ with $\phi=4$.
We train the neural network using $R=250$ posterior samples by applying double bps to each generated dataset with $K=5$ subsets, $\alpha\in\{0.7,0.8,0.9\}$ and $\phi\in\{3,4,5\}$. These yield $100$ instances of $\{Z,\Theta\}$, where $Z=[Y:X]\in\mathbb{R}^{n\times(q+p)}$ and $\Theta\in\mathbb{R}^{[(qp)+(q(q+1)/2)+(nq)]\times 3}$ comprises the $\{2.5,50,97.5\}$ posterior quantiles for the distinct elements of $\{\beta,\Sigma,\Omega\}$.
We use a deep neural network comprising 3 hidden layers with 128, 256, and 512 nodes, with ReLU activations. The residual network is trained over 50 epochs with 24 batches per epoch. For evaluation, we apply the trained model to unseen datasets with the same dimensions. Figure 4 displays the predicted posterior intervals (blue bars) for $\{\beta,\Sigma\}$, along with the true values of the parameters (yellow triangles). The credible intervals $95\%$ capture all true values and the predicted medians ($50\%$) align closely with the true values, highlighting the strong predictive precision of the deep network in recovering posterior summaries. Figure 5 displays the predicted posteriors for $\Omega$ quantiles. We compare the results from amortized inference, with the true $\Omega$, and the double bps prediction for $50$th quantile, presented in the first and the second column of Figure 5, respectively.
These results illustrate the strengths of amortized inference and transfer learning. Once trained, the deep network provides instant posterior quantile estimates for new datasets, without requiring us to rerun double bps. This amortizes the computational cost for future tasks. Additionally, the model generalizes across a range of data-generating conditions effectively enabling posterior transfer learning to new but structurally similar problems. This makes the approach especially useful in large-scale or resource-constrained applications where repeated full Bayesian inference is prohibitively expensive.